OpenStudy (anonymous):

find the equation of the line passing through (0,-1),(1,3). Enter the equation as y=mx+b.

OpenStudy (anonymous):

f(x)=4x-1

hero (hero):

Steps should be explained, not just answers given

OpenStudy (anonymous):

are u sure? I put y=4x-1 into graph and it doesn't pass thru those points? because thats what i got too

hero (hero):

@Kathi26

OpenStudy (anonymous):

@Hero is that the answer? i got y=4x-1

hero (hero):

Let me see

OpenStudy (anonymous):

ok. after, can u help me with a few more? then thats it? since I see you're smart:)

hero (hero):

You probably won't understand those steps

hero (hero):

There's more than one way to do these

OpenStudy (anonymous):

So it is not what the other person said, its y=2x-1?

hero (hero):

Hang on a sec

OpenStudy (anonymous):

but the slope is 4..

hero (hero):

I know, I goofed up, that's why I said hang on a sec

hero (hero):

(x,y) = (0,-1) y = mx + b -1 = m(0) +b -1 = b (x,y) = (1,3) y = mx + b 3 = m(1) - 1 3+1 = m 4 = m y = 4x - 1

hero (hero):

So that's the steps and 4 is the slope

OpenStudy (anonymous):

Ok can u hlp with a few more! thanks so muuhc: which of these points lie on the line that passes through (6,2) with slope m=2/3? Select all answers that are correct. A) 3,0 B) (-4,0) C) (3,-2) D) (-3,-4)

hero (hero):

You have to use the point slope formula \((y - y_1) = m(x - x_1)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for the first ordered pair, i got y-0=2/3(x-0)

hero (hero):

The first ordered pair is not (0,0)

OpenStudy (anonymous):

im not sure how to check if the ordered pairs work

OpenStudy (anonymous):

and i meant x-3

hero (hero):

Well, you're supposed to use the point-slope formula to get the form y = mx+b

hero (hero):

But you don't really have to do it that way

hero (hero):

You can just continue using y = mx + b and substituting values to find m and b

OpenStudy (anonymous):

Can u do that? i tried and can't get it

hero (hero):

Basically, they give you one of the points (6,2) so we know y = mx + b becomes: 2 = 2/3(6) + b

hero (hero):

Now just solve for b

OpenStudy (anonymous):

Ok so b is -2

hero (hero):

Very good

hero (hero):

Now you have y = 2/3(x) - 2

hero (hero):

What you need to do now is substitute the other points and see if you get equal numbers on both sides of the equal sign

OpenStudy (anonymous):

then plug in ordered pairs. let me do that and tell u what i gott

OpenStudy (anonymous):

I got A and D

OpenStudy (anonymous):

3,0 and -3,-4

hero (hero):

Yup, you got it :D